I am having trouble converting this to an equation. I know the answer but can't come up with the equation that will give me the correct answer.
There is a group of boys and dogs. There are 22 heads and 68 feet. How many boys are there and how many dogs are there.
The answer is 10 boys and 12 dogs,but I need an equation.
Thank You

Answer by checkley71(8403)

HEADS=B+D
22=B+D
B=22-D
FEET=2B+4D
68=2B+4D SUBSTITUTE (22-D) FOR B IN THE FEET EQUATION WE GET
68=2(22-D)+4D
68=44-2D+4D
68-44=2D
24=2D
D=24/2
D=12 ANSWER
22=B+12
B=22-12
B=10 ANSWER
PROOF
68=2*10+4*12
68=20+48
68=68